# Motives of nullcones of quiver representations

### v2

@@ 3. Applications and examples
    It is left as an exercise to the reader that, for general $d$, a monomial $\mathbb{L}^{a}$ appears in $[N_{d}]$ with non-zero coefficient if and only if $a=km+l$ for $d-1\leq k\leq{d\choose 2}$ and (${d\choose 2}-k\leq l\leq{d\choose 2}$.
4. Hesselink stratification and rational/integral level quiver
4.1. Introduction
@@
We want to calculate the motive of $N_{\mathbf{d}}(Q)$ as the corresponding element in the Grothendieck ring of varieties using the Hesselink stratification of the nullcone. To describe the Hesselink stratification of $N_{\mathbf{d}}(Q)$, we interpret the results of [8].
@@
The Hesselink stratification decomposes the nullcone into finitely many locally closed subsets $S_{s}$, so in the Grothendieck ring of varieties we obtain the identity of motives:
(1) $[N_{\mathbf{d}}]=\sum_{s\in L_{\mathbf{d}}}[S_{s}]~{},$
@@
where $L_{\mathbf{d}}$ denotes the index set of strata. Our aim is to write the set of strata in terms of so-called dominant and balanced strings. This description covers all strata, albeit including empty ones. Depending on $s$ we construct a tuple $(Q_{s},\theta_{s},{\bf d}_{s})$ consisting of a quiver, a stability and a dimension vector, such that the stratum $S_{s}$ is non-empty if and only if there exists a $\theta_{s}$-semistable representations in $R_{{\bf d}_{s}}(Q_{s})$.
@@
Starting with a quiver $Q$ with $|Q_{0}|=n$ vertices, set of arrows $Q_{1}$, and a dimension vector $\mathbf{d}=(d_{i})_{i}$, we denote $d=\sum_{i=1}^{n}d_{i}$. We decompose the interval
@@
We fix a maximal torus $T_{\mathbf{d}}$ of $G_{\mathbf{d}}$ and denote by $\Pi$ the set of weights of $T_{\mathbf{d}}$ having a nontrivial weight space in $R_{\mathbf{d}}(Q)$. Explicitly, the group of weights of $T_{\mathbf{d}}$ is isomorphic to $\mathbb{Z}^{d}$ via the canonical generators $\pi_{i}$ for $1\leq i\leq d$. Denoting $\pi_{ij}=\pi_{j}-\pi_{i}$ we obtain
$\Pi=\{\pi_{ij}=\pi_{j}-\pi_{i}\;|\;i\in I_{v},j\in I_{v^{\prime}}\text{ for some arrow }\alpha:v\rightarrow v^{\prime}\text{ in }Q\}~{}.$
@@
a) A subset $R\subset\Pi$ is said to be unstable if there exists a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ such that $s_{j}-s_{i}\geq 1$ for all $\pi_{ij}\in R$.
If $R$ is unstable, there is a unique coweight $s(R)\in\mathbb{Q}^{d}$ (see [9, Lemma 6.2]) with this property and such that the norm $|s(R)|=s_{1}^{2}+\dots+s_{d}^{2}$ is minimal. We will call such a coweight minimal.
b) For a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ we define the corresponding subset $R(s)\subset\Pi$ as
$R(s)=\{\pi_{ij}\in\Pi\;|\;s_{j}-s_{i}\geq 1\}$
c) We call $R$ a saturated subset of $\Pi$ whenever $R=R(s(R))$.
d) We denote by $S_{\mathbf{d}}=S_{d_{1}}\times\dots\times S_{d_{n}}$ the Weyl group of $G_{\mathbf{d}}$.
Remark 4.2.
(see [8, Section 2.2]) If $R$ is a saturated subset of $\Pi$, we have a corresponding saturated subspace $X_{R}$ of $N_{\mathbf{d}}(Q)$ by taking
$X_{R}=\bigoplus_{\pi_{ij}\in R}R_{\mathbf{d}}(Q)_{\pi_{ij}}~{}.$
By [5] there is a bijection between the $G_{\mathbf{d}}$-conjugacy classes of saturated subspaces of $N_{\mathbf{d}}(Q)$ and the conjugacy classes of saturated subsets $R\subset\Pi$ under the action of the Weyl group $S_{\mathbf{d}}$. This correspondence assigns to $R$ the subspace $X_{R}$, and to a saturated subspace the set of nonzero weights of its elements. Hence, the number of Hesselink strata of $N_{\mathbf{d}}(Q)$ is at most the number of conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group.
We construct a bijection between the conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group and coweights satisfying certain additional properties. Consider a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ for $G_{\mathbf{d}}$. Then there is a unique partition of $[1\ldots d]$ into a minimal number of pairwise disjoint subsets $\tilde{J}$, such that the $s_{j}$ for $j\in\tilde{J}$ form an ordered sequence increasing by 00 or $1$ in each step after sorting. We call such a sequence a segment and denote it by $J$. After introducing this notation we can now state definitions for the required conditions on coweights.
@@
a) A coweight is called balanced if and only if for every segment $J$ of $s$ we have
$\sum_{i\in\tilde{J}}s_{i}=0~{}.$
b) We call a coweight $s\in\mathbb{Q}^{d}$ dominant if and only if for every vertex $v$ we have, for all $i,j\in I_{v}$,
$s_{i}\leq s_{j}\text{ whenever }i\leq j~{}.$
c) We denote by $\tilde{L}_{\mathbf{d}}$ the set of dominant and minimal coweights.
Since every saturated subset $R$ is determined by its unique minimal coweight $s(R)$, and every minimal coweight belongs to exactly one saturated subset, we have a bijection between the conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group and minimal dominant coweights.
Since the property of being minimal can be hard to verify, our goal is to extend the index set in such a way, that the new summands in the equation (1) are equal to zero, and it is easy to verify if a string belongs to the new index set. The property of being balanced will play an essential role and the following lemma portrays the first step [8], see also [5].
@@
Suppose $\mu=(\mu_{1},\dots,\mu_{d})$ is not balanced. Then there exists a segment $J$ such that $\sum_{i\in\tilde{J}}\mu_{i}\neq 0$. By definition of the segments, for every $\mu_{j}\in J$ and $\mu_{k}\notin J$ we have $|\mu_{j}-\mu_{k}|<1$ or $|\mu_{j}-\mu_{k}|>1$. Consequently we can choose an $\epsilon>0$ such that these inequalities also hold for $\mu_{j}+\epsilon$ and $\mu_{j}-\epsilon$. By shifting all entries of $J$ by $\epsilon$ or $-\epsilon$, the distance between two entries in the segment does not change. In this way we construct another string $\mu^{\prime}$ such that $R(\mu)=R(\mu^{\prime})$. Without loss of generality let $\sum_{i\in\tilde{J}}\mu_{i}=\alpha>0$ and we define $\mu^{\prime}$ by
$\mu_{i}^{\prime}=\begin{cases}\mu_{i}&\text{if }i\notin\tilde{J}\\ \mu_{i}-\epsilon&\text{if }i\in\tilde{J}~{}.\end{cases}$
We denote the cardinality of $\tilde{J}$ by $n$ and obtain:
@@
$=\sum_{\tilde{J}}-2\mu_{i}\epsilon+\epsilon^{2}=-2\alpha\epsilon+n\epsilon^{2}~{}.$
@@
$\mu_{i}^{\prime}=\begin{cases}\mu_{i}&\text{if }i\notin\tilde{J}\\ \mu_{i}+\epsilon&\text{if }i\in\tilde{J}~{}.\end{cases}$
We can thus extend the index set of the Hesselink strata using balanced and dominant coweights. The problem of detecting which strata are non-empty remains. To this effect, for every coweight $s$, we define a tuple $(Q_{s},{\bf d}_{s},\theta_{s})$, consisting of a quiver, a dimension vector and a stability, as follows:
@@
For each segment $J_{i}$, where $1\leq i\leq u$, we consider the level quiver $Q_{i}$ on $n\times(k_{i}+1)$ vertices
$\{(v,j)\;|\;v\in Q_{0}\text{ and }0\leq j\leq k_{i}\}~{},$
where $p_{i}+k_{i}$ is the last entry in $J_{i}$. In $Q_{i}$ the number of arrows from $(v,k)$ to $(v^{\prime},k+1)$ equals the number of arrows from $v$ to $v^{\prime}$ in $Q$. The dimension vector ${\bf d}_{i}$ of $Q_{i}$ is given by $(a_{i,k}^{v})_{v,k}$.
@@
A stability on $Q_{s}$ is defined by $\theta_{s}=(m_{v,j}^{(i)})_{i,v,j}$, where $m_{v,j}^{(i)}=-z\cdot(p_{i}+j)$ and $z$ is the least common multiple of all denominators of the rational numbers $p_{i}$ with $1\leq i\leq u$.
We recall that a representation $V$ of a quiver $Q$ of dimension vector ${\bf d}$ is called $\theta$-semistable if $\theta({\bf d})=0$ and $\theta({\bf e})\leq 0$ whenever ${\bf e}$ is the dimension vector of a subrepresentation of $V$ (note that this definition uses the opposite sign convention to [8]). We then denote by $R_{\bf d}^{\theta-sst}(Q)=R_{\bf d}^{sst}(Q)$ the $\theta$-semistable locus in $R_{\bf d}(Q)$. More generally, using the slope function $\sigma=\theta/\dim$ (and not requiring $\theta({\bf d})=0$) we call the representation $V$ $\sigma$-semistable if $\sigma({\bf e})\leq\sigma({\bf d})$ for all dimension vectors ${\bf e}$ of non-zero subrepresentations of $V$.
The following theorem of [8] summarizes the main properties relating the representation spaces of level quivers and the Hesselink stratification:
@@
The elements $s$ of $\tilde{L}_{\mathbf{d}}$ give rise to Hesselink strata $S_{s}$, where $S_{s}$ is non-empty if and only if the set of $\theta_{s}$-semistable representations in $R_{{\bf d}_{s}}(Q_{s})$ is non-empty. The Hesselink strata are isomorphic to associated fibre bundles
$S_{s}\cong G_{\mathbf{d}}\times^{P_{s}}V_{s}^{sst}~{}.$
Here $P_{s}$ is a certain parabolic subgroup of $G_{\mathbf{d}}$ and $V_{s}^{sst}$ is given as follows:
There exists a vector bundle
$p_{s}:V_{s}\rightarrow R_{{\bf d}_{s}}(Q_{s})~{},$
such that
$V_{s}^{sst}=p_{s}^{-1}(R_{{\bf d}_{s}}^{\theta_{s}-sst}(Q_{s}))~{}.$
In particular, $P_{s},V_{s}$ and $R_{{\bf d}_{s}}(Q_{s})$ can be described directly in terms of $s\in\tilde{L}_{\mathbf{d}}$, namely:
@@
Consider $|Q_{1}|$-tuples of $d\times d$-matrices. Then $V_{s}$ is given as the affine space of such tuples, where the matrix corresponding to an arrow $\alpha:v\rightarrow v^{\prime}$ has arbitrary entries at a position $(j,i)$ if $i\in I_{v}$, $j\in I_{v^{\prime}}$ and $s_{j}-s_{i}\geq 1$, and zero entries everywhere else (compare with Definition 4.1). The group $P_{s}$ is given by a single invertible $d\times d$-matrix with zero entries everywhere except possibly at a position $(j,i)$ if $i,j\in I_{v}$ for some vertex $v\in Q_{0}$ and $s_{j}-s_{i}\geq 0$. For the affine subspace $B_{s}\subset V_{s}$ obtained by additionally requiring zero entries at all positions $(j,i)$ except the ones where $s_{j}-s_{i}=1$, we have
$B_{s}\cong R_{{\bf d}_{s}}(Q_{s}).$
@@
Now we can again consider the identity of motives in the localized base ring $R$ from the beginning:
$[N_{\mathbf{d}}]=\sum_{s\in L_{\mathbf{d}}}[S_{s}]~{}.$
@@
We divide by the motive of $G_{\mathbf{d}}$ and use the isomorphism $S_{s}\cong G_{\mathbf{d}}\times^{P_{s}}V_{s}^{sst}$ to obtain the following:
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}$ $=\sum_{s\in L_{\mathbf{d}}}\frac{[S_{s}]}{[G_{\mathbf{d}}]}=\sum_{s\in\tilde{L}_{\mathbf{d}}}\frac{[G_{\mathbf{d}}\times^{P_{s}}V_{s}^{sst}]}{[G_{\mathbf{d}}]}~{}.$
@@
Since the group $P_{s}$ is special we know that
$[G_{\mathbf{d}}\times^{P_{s}}V_{s}^{sst}]\cdot[P_{s}]=[G_{\mathbf{d}}]\cdot[V_{s}^{sst}]$
and find
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in\tilde{L}_{\mathbf{d}}}\frac{[V_{s}^{sst}]}{[P_{s}]}~{}.$
Recall from Theorem 4.6 that there is a vector bundle
$p_{s}:V_{s}\rightarrow R_{\mathbf{d}_{s}}(Q_{s})$
of rank $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$. Since $R^{sst}_{\mathbf{d}_{s}}(Q_{s})$ is Zariski-open in $R_{\mathbf{d}_{s}}(Q_{s})$ we also obtain a vector bundle
$p^{\prime}_{s}:V_{s}^{sst}\rightarrow R^{sst}_{\mathbf{d}_{s}}(Q_{s})$
of the same rank as long as $R^{sst}_{\mathbf{d}_{s}}(Q_{s})$ is not empty. Moreover, the reductive part of $P_{s}$ is isomorphic to $G_{\mathbf{d}_{s}}$, yielding
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in\tilde{L}_{\mathbf{d}}}\frac{[R^{sst}_{\mathbf{d}_{s}}(Q_{s})]\cdot\mathbb{L}^{\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})}}{[G_{\mathbf{d}_{s}}]\cdot\mathbb{L}^{\dim P_{s}-\dim G_{\mathbf{d}_{s}}}}~{}.$

### v3

@@ 3.2 Examples
    It is left as an exercise to the reader that, for general $d$, a monomial $\mathbb{L}^{a}$ appears in $N_{d}]$ with non-zero coefficient if and only if $a=km+l$ for $d-1\leq k\leq{d\choose 2}$ and ${d\choose 2}-k\leq l\leq{d\choose 2}$.
  5. (5)
    The proof of [1, Lemma 2.6] can be read as a motivic identity, resulting in the following formula of a rather different nature:
    $N_{\bf d}]/[G_{\bf d}]=\sum_{({\bf d}^{k})_{k}}\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle-\sum_{k}\sum_{i\in Q_{0}}(d^{k}_{i})^{2}}\prod_{k}\prod_{i\in Q_{0}}\frac{\bigl(\mathbb{L}^{-1}\bigr)_{\sum_{\alpha\colon j\rightarrow i}d^{k}_{j}}}{\bigl(\mathbb{L}^{-1}\bigr)_{d^{k}_{i}}\bigl(\mathbb{L}^{-1}\bigr)_{\sum_{\alpha\colon j\rightarrow i}d^{k}_{j}-d^{k+1}_{i}}},$
    where the sum ranges over all ordered decompositions ${\bf d}=\sum_{k}{\bf d}^{k}$.
4 Hesselink stratification and rational/integral level quiver
4.1 Introduction
@@
We want to calculate the motive of $N_{\mathbf{d}}$ as the corresponding element in the Grothendieck ring of varieties using the Hesselink stratification of the nullcone. To describe the Hesselink stratification of $N_{\mathbf{d}}$, we interpret in detail the results of [8], thereby setting up notation. The first new results will only be derived after Proposition 4.8.
In fact, we will never use (and consequently, never state) the actual definition of these strata themselves, which is rather indirect (very briefly, one considers the $G_{\bf d}$-saturations of equivalence classes of points in $N_{\bf d}$ under the equivalence relation that their sets of so-called optimal one-parameter subgroups coincide). What is only needed for describing the motive of $N_{\bf d}$ is the isomorphism type of the Hesselink strata, which is described in Theorem 4.7, and an explicit parametrization of the strata.
So we will mainly be concerned with the index set of Hesselink strata $L_{\bf d}$, which we will adapt several times. We will provide an overview over the next steps for more convenience and lucidity. Originally, $L_{\bf d}$ is defined in terms of so-called saturated sets of weights in $R_{\bf d}(Q)$, or, equivalently, in terms of certain so-called dominant minimal coweights $s$. Our first goal is to extend the index set $L_{\mathbf{d}}$ to the set of all dominant and minimal coweights $\widetilde{L}_{\mathbf{d}}$, see Remark 4.2 and the subsequent paragraphs. In a next step we use Theorem 4.12 to extend $\widetilde{L}_{\mathbf{d}}$ to the set of dominant and balanced coweights $\overline{L}_{\mathbf{d}}$. Afterwards we show that $\overline{L}_{\mathbf{d}}$ is in bijection to $L^{\prime}_{\mathbf{d}}$, the set of balanced dimension vectors for the so-called rational level quiver $Q_{\mathbb{Q}}$ (see Proposition 4.16) and ultimately we prove that $L^{\prime}_{\mathbf{d}}$ is in bijection to $L(\mathbf{d})$, a set of tuples of dimension vectors for the so-called integral level quiver, satisfying certain additional requirements (see Proposition 4.17). We summarize and finish this overview with a diagram of index sets
$L_{\bf d}\subset\widetilde{L}_{\bf d}\subset\overline{L}_{\bf d}\leftrightarrow L_{\bf d}^{\prime}\leftrightarrow L({\bf d}).$
@@
Since the Hesselink stratification decomposes the nullcone into finitely many locally closed subsets $S_{s}$, we obtain the following identity of motives in the Grothendieck ring of varieties:
$N_{\mathbf{d}}]=\sum_{s\in L_{\mathbf{d}}}[S_{s}].$ (4.1)
@@
Our first goal is to write the set of strata $L_{\bf d}$ in terms of so-called dominant and balanced strings. This description covers all strata, albeit including empty ones. Depending on $s$ we construct a tuple $(Q_{s},\theta_{s},{\bf d}_{s})$ consisting of a quiver, a stability and a dimension vector, such that the stratum $S_{s}$ is non-empty if and only if there exists a $\theta_{s}$-semistable representations in $R_{{\bf d}_{s}}(Q_{s})$.
@@
Starting with a quiver $Q$ with $|Q_{0}|=n$ vertices, set of arrows $Q_{1}$, and a dimension vector $\mathbf{d}=(d_{i})_{i}$, we denote $d=\dim{\bf d}$. We decompose the interval
@@
We fix a maximal torus $T_{\mathbf{d}}$ of $G_{\mathbf{d}}$ and denote by $\Pi$ the set of weights of $T_{\mathbf{d}}$ having a nontrivial weight space in $R_{\mathbf{d}}(Q)$. Explicitly, the group of weights of $T_{\mathbf{d}}$ is isomorphic to $\mathbb{Z}^{d}$ via the canonical generators $\pi_{i}$ for $1\leq i\leq d$. Denoting $\pi_{ij}=\pi_{j}-\pi_{i}$, we obtain
$\Pi=\{\pi_{ij}=\pi_{j}-\pi_{i}\mid i\in I_{v},\,j\in I_{v^{\prime}}\text{ for some arrow }\alpha\colon v\rightarrow v^{\prime}\text{ in }Q\}$
and we denote by $R_{\mathbf{d}}(Q)_{\pi_{ij}}$ the corresponding weight space.
@@
  • (a)
    A subset $\mathcal{R}\subset\Pi$ is said to be unstable if there exists a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ such that $s_{j}-s_{i}\geq 1$ for all $\pi_{ij}\in\mathcal{R}$. If $\mathcal{R}$ is unstable, there is a unique coweight $s(\mathcal{R})\in\mathbb{Q}^{d}$ (see [9, Lemma 6.2]) with this property and such that the norm(-square) $|s(\mathcal{R})|=s_{1}^{2}+\dots+s_{d}^{2}$ is minimal. We will call such a coweight minimal.
  • (b)
    For a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ we define the corresponding subset $\mathcal{R}(s)\subset\Pi$ as
    $\mathcal{R}(s)=\{\pi_{ij}\in\Pi\mid s_{j}-s_{i}\geq 1\}.$
  • (c)
    We call $\mathcal{R}$ a saturated subset of $\Pi$ whenever $\mathcal{R}=\mathcal{R}(s(\mathcal{R}))$.
  • (d)
    We denote by $S_{\mathbf{d}}=S_{d_{1}}\times\dots\times S_{d_{n}}$ the Weyl group of $G_{\mathbf{d}}$.
Using this terminology, we can now describe the indexing set for the Hesselink strata; the formal definition will follow in Definition 4.3.
Remark 4.2 (see [8, Section 2.2]).
If $\mathcal{R}$ is a saturated subset of $\Pi$, we have a corresponding saturated subspace $X_{\mathcal{R}}$ of $N_{\mathbf{d}}$ by taking
$X_{\mathcal{R}}=\bigoplus_{\pi_{ij}\in\mathcal{R}}R_{\mathbf{d}}(Q)_{\pi_{ij}}.$
By [8, Section 2.2], every Hesselink stratum is a Zariski-open subset of some $X_{\mathcal{R}}$. We can thus label the Hesselink strata by certain saturated sets of weights $\mathcal{R}$, determined up to conjugacy. Namely, by [5] there is a bijection between the $G_{\mathbf{d}}$-conjugacy classes of saturated subspaces of $N_{\mathbf{d}}$ and the conjugacy classes of saturated subsets $\mathcal{R}\subset\Pi$ under the action of the Weyl group $S_{\mathbf{d}}$. This correspondence assigns to $\mathcal{R}$ the subspace $X_{\mathcal{R}}$, and to a saturated subspace the set of nonzero weights of its elements. In particular, the number of Hesselink strata of $N_{\mathbf{d}}$ is at most the number of conjugacy classes of saturated subsets $\mathcal{R}\subset\Pi$ under the Weyl group.
We construct a bijection between the conjugacy classes of saturated subsets $\mathcal{R}\subset\Pi$ under the Weyl group and coweights satisfying certain additional properties. Consider a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ for $G_{\mathbf{d}}$. Then there is a unique partition of $1\dots d]$ into a minimal number of pairwise disjoint subsets $\tilde{J}$, such that the $s_{j}$ for $j\in\tilde{J}$ form an ordered sequence increasing by 0 or $1$ in each step after sorting. We call such a sequence a segment and denote it by $J$.
A clarifying example: Consider the coweight
$\left(\frac{3}{2},0,-\frac{1}{2},\frac{4}{3},1,-\frac{1}{2},\frac{1}{3},\frac{1}{2},-1,\frac{3}{2}\right)\in\mathbb{Q}^{10}.$
The corresponding segments are
$J_{1}$ $=\left(-\frac{1}{2},-\frac{1}{2},\frac{1}{2},\frac{3}{2},\frac{3}{2}\right),\qquad J_{2}=(-1,0,1),\qquad J_{3}=\left(\frac{1}{3},\frac{4}{3}\right)$
and the corresponding partition of $1\dots 10]$ is $\tilde{J}_{1}=[3,6,8,1,10]$, $\tilde{J}_{2}=[9,2,5]$ and $\tilde{J}_{3}=[7,4]$.
After introducing this notation, we can now state definitions for the required conditions on coweights.
@@
  • (a)
    A coweight is called balanced if and only if for every segment $J$ of $s$ we have $\sum_{i\in\tilde{J}}s_{i}=0$.
  • (b)
    We call a coweight $s\in\mathbb{Q}^{d}$ dominant if and only if for every vertex $v$ we have, for all $i,j\in I_{v}$, $s_{i}\leq s_{j}$ whenever $i\leq j$.
  • (c)
    We denote by $L_{\bf d}$ the set of dominant minimal coweights for which $X_{\mathcal{R}(s)}$ is the closure of a Hesselink stratum.
  • (d)
    We denote by $\widetilde{L}_{\mathbf{d}}$ the set of all dominant and minimal coweights.
Since every saturated subset $\mathcal{R}$ is determined by its unique minimal coweight $s(\mathcal{R})$, and every minimal coweight belongs to exactly one saturated subset, we have a bijection between the conjugacy classes of saturated subsets $\mathcal{R}\subset\Pi$ under the Weyl group and minimal dominant coweights.
Since the property of being minimal can be hard to verify, our goal is to extend the index set $\widetilde{L}_{\mathbf{d}}$ to $\overline{L}_{\mathbf{d}}$ in such a way, that the new summands in the equation (4.1) are equal to zero, and such that it is easy to verify if a string belongs to the new index set. The property of being balanced will play an essential role in extending the index set of the Hesselink strata. The following lemma portrays the first step [8], see also [5].
@@
Suppose $\mu=(\mu_{1},\dots,\mu_{d})$ is not balanced. Then there exists a segment $J$ such that $\sum_{i\in\tilde{J}}\mu_{i}\neq 0$. By definition of the segments, for every $\mu_{j}\in J$ and $\mu_{k}\notin J$ we have $|\mu_{j}-\mu_{k}|<1$ or $|\mu_{j}-\mu_{k}|>1$. Consequently, we can choose an $\epsilon>0$ such that these inequalities also hold for $\mu_{j}+\epsilon$} and $\mu_{j}-\epsilon$. By shifting all entries of $J$ by $\epsilon$ or $\epsilon$, the distance between two entries in the segment does not change. In this way, we construct another string $\mu^{\prime}$ such that $\mathcal{R}(\mu)=\mathcal{R}(\mu^{\prime})$. First assuming $\sum_{i\in\tilde{J}}\mu_{i}=\alpha>0$ and we define $\mu^{\prime}$ by
[…]
We denote the cardinality of $\tilde{J}$ by $n$ and obtain
@@
$=\sum_{\tilde{J}}-2\mu_{i}\epsilon+\epsilon^{2}=-2\alpha\epsilon+n\epsilon^{2}.$
@@
[…]
We can thus extend the index set of the Hesselink strata using balanced and dominant coweights (this will be done before Lemma 4.9). The problem of detecting which strata are non-empty remains. To this effect, for every coweight $s$, we define a tuple $(Q_{s},{\bf d}_{s},\theta_{s})$, consisting of a quiver, a dimension vector and a stability, as follows.
@@
For each segment $J_{i}$, where $1\leq i\leq u$, we consider the level quiver $Q_{i}$ on $n\times(k_{i}+1)$ vertices $\{(v,j)\mid v\in Q_{0}\text{ and }0\leq j\leq k_{i}\}$}, where $p_{i}+k_{i}$ is the last entry in $J_{i}$. In $Q_{i}$ the number of arrows from $(v,k)$ to $(v^{\prime},k+1)$ equals the number of arrows from $v$ to $v^{\prime}$ in $Q$. The dimension vector ${\bf d}_{i}$ of $Q_{i}$ is given by $(a_{i,k}^{v})_{v,k}$.
Definition 4.6.
@@
A stability on $Q_{s}$ is defined by $\theta_{s}=\bigl(m_{v,j}^{(i)}\bigr)_{i,v,j}$, where $m_{v,j}^{(i)}=-z(p_{i}+j)$ and $z$ is the least common multiple of all denominators of the rational numbers $p_{i}$ with $1\leq i\leq u$. In fact, the factor $z$ is to ensure compatibility with [8].
We recall that a representation $V$ of a quiver $Q$ of dimension vector $\bf d$} is called $\theta$-semistable if $\theta({\bf d})=0$} and $\theta({\bf e})\leq 0$ whenever $\bf e$} is the dimension vector of a subrepresentation of $V$ (note that this definition uses the opposite sign convention to [8]). We then denote by $R_{\bf d}^{\theta\text{\rm-sst}}(Q)=R_{\bf d}^{\mathrm{sst}}(Q)$ the $\theta$-semistable locus in $R_{\bf d}(Q)$. More generally, using the slope function $\sigma=\theta/{\dim}$ (and not requiring $\theta({\bf d})=0$) we call the representation $V$ $\sigma$-semistable if $\sigma({\bf e})\leq\sigma({\bf d})$ for all dimension vectors $\bf e$} of non-zero subrepresentations of $V$.
The following theorem of [8] summarizes the main properties relating the representation spaces of level quivers and the Hesselink stratification.
@@
The elements $s$ of $\widetilde{L}_{\mathbf{d}}$ give rise to Hesselink strata $S_{s}$, where $S_{s}$ is non-empty if and only if the set of $\theta_{s}$-semistable representations in $R_{{\bf d}_{s}}(Q_{s})$ is non-empty. The Hesselink strata are isomorphic to associated fibre bundles $S_{s}\cong G_{\mathbf{d}}\times^{P_{s}}V_{s}^{\mathrm{sst}}$. Here $P_{s}$ is a certain parabolic subgroup of $G_{\mathbf{d}}$ and $V_{s}^{\mathrm{sst}}$ is given as follows:
There exists a vector bundle $p_{s}\colon V_{s}\rightarrow R_{{\bf d}_{s}}(Q_{s})$, such that $V_{s}^{\mathrm{sst}}=p_{s}^{-1}\bigl(R_{{\bf d}_{s}}^{\theta_{s}\text{\rm-sst}}(Q_{s})\bigr)$.
Proposition 4.8.
@@
$P_{s}$, $V_{s}$ and $R_{{\bf d}_{s}}(Q_{s})$ can be described directly in terms of $s\in\widetilde{L}_{\mathbf{d}}$, namely: Consider $|Q_{1}|$-tuples of $d\times d$ matrices. Then $V_{s}$ is given as the affine space of such tuples, where the matrix corresponding to an arrow $\alpha\colon v\rightarrow v^{\prime}$ has arbitrary entries at a position $(j,i)$ if $i\in I_{v}$}, $j\in I_{v^{\prime}}$ and $s_{j}-s_{i}\geq 1$, and zero entries everywhere else ((compare with Definition 4.1). The group $P_{s}$ is given by a single invertible $d\times d$ matrix with zero entries everywhere except possibly at a position $(j,i)$ if $i,j\in I_{v}$ for some vertex $v\in Q_{0}$ and $s_{j}-s_{i}\geq 0$. For the affine subspace $B_{s}\subset V_{s}$ obtained by additionally requiring zero entries at all positions $(j,i)$ except the ones where $s_{j}-s_{i}=1$, we have $B_{s}\cong R_{{\bf d}_{s}}(Q_{s})$.
@@
Now we can again consider the identity of motives in the localized base ring $\mathbb{M}$ from the beginning of this section
$[N_{\mathbf{d}}]=\sum_{s\in L_{\mathbf{d}}}[S_{s}].$
@@
We divide by the motive of $G_{\mathbf{d}}$ and use the isomorphism $S_{s}\cong G_{\mathbf{d}}\times^{P_{s}}V_{s}^{\mathrm{sst}}$ to obtain the following:
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}$ $=\sum_{s\in L_{\mathbf{d}}}\frac{[S_{s}]}{[G_{\mathbf{d}}]}=\sum_{s\in\widetilde{L}_{\mathbf{d}}}\frac{\big[G_{\mathbf{d}}\times^{P_{s}}V_{s}^{\mathrm{sst}}\big]}{[G_{\mathbf{d}}]}.$
@@
Since the group $P_{s}$ is special, we know that $\big[G_{\mathbf{d}}\times^{P_{s}}V_{s}^{\mathrm{sst}}\big]\cdot[P_{s}]=[G_{\mathbf{d}}]\cdot\big[V_{s}^{\mathrm{sst}}\big]$ and find
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in\widetilde{L}_{\mathbf{d}}}\frac{\big[V_{s}^{\mathrm{sst}}\big]}{[P_{s}]}.$
Recall from Theorem 4.7 that there is a vector bundle $p_{s}\colon V_{s}\rightarrow R_{\mathbf{d}_{s}}(Q_{s})$ of rank $\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})$. Since $R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})$ is Zariski-open in $R_{\mathbf{d}_{s}}(Q_{s})$ we also obtain a vector bundle $p^{\prime}_{s}\colon V_{s}^{\mathrm{sst}}\allowbreak\rightarrow R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})$ of the same rank as long as $R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})$ is not empty. Moreover, the reductive part of $P_{s}$ is isomorphic to $G_{\mathbf{d}_{s}}$, yielding
$\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}=\sum_{s\in\widetilde{L}_{\mathbf{d}}}\frac{\big[R^{\mathrm{sst}}_{\mathbf{d}_{s}}(Q_{s})\big]\cdot\mathbb{L}^{\dim V_{s}-\dim R_{\mathbf{d}_{s}}(Q_{s})}}{[G_{\mathbf{d}_{s}}]\cdot\mathbb{L}^{\dim P_{s}-\dim G_{\mathbf{d}_{s}}}}.$